### v1

# Lifting the Veil on the Large Language Model Supply Chain: Composition, Risks, and Mitigations

Abstract.
Large language models (LLM) have sparked significant impact with regard to both intelligence and productivity. In recent years, a great surge has been witnessed in the introduction of both commercial and open-source LLMs. Many businesses have adopted the LLMs into their applications to solve their own domain-specific tasks. However, integrating LLMs into specific business scenarios requires more than just utilizing the models themselves. Instead, it is a systematic process that involves substantial components, which are collectively referred to as the LLM supply chain. The LLM supply chain inherently carries risks. Therefore, it is essential to understand the types of components that may be introduced into the supply chain and the associated risks, enabling different stakeholders to implement effective mitigation measures. While some literature touches on risks associated with the LLM supply chain, there is currently no paper that explicitly defines its scope, identifies inherent risks, and examines potential mitigation strategies. As LLMs have become essential infrastructure in the new era, we believe that a thorough review of the LLM supply chain, along with its inherent risks and mitigation strategies, would be valuable for industry practitioners to avoid potential damages and losses, and enlightening for academic researchers to rethink existing approaches and explore new avenues of research. Our paper provides a comprehensive overview of the LLM supply chain, detailing the stakeholders, composing artifacts, and the supplying types. We developed taxonomies of risk types, risky actions, and mitigations related to various supply chain stakeholders and components. In summary, our work explores the technical and operational aspects of the LLM supply chain, offering valuable insights for researchers and engineers in the evolving LLM landscape.
1. Introduction
The large language models (LLMs) have sparked unprecedented discussion about the power of generative language models, showcasing their remarkable ability to understand human-written text, embody vast knowledge, and generate responses based on user instructions. In 2023 following ChatGPT’s debut, there has been a significant surge in the development and introduction of large language models, regarding both commercial LLMs and free open-source LLMs. Researchers and practitioners have extended the capabilities of LLMs beyond natural language processing, applying them in fields like software engineering, finance, and education, dramatically reshaping these areas (Science, 2024; Fan et al., 2023; bloomberg, 2024).
Integrating LLMs involves more than just utilizing the models themselves. On one hand, although some applications may directly employ commercial LLMs, they still require extra solutions for pre-processing (e.g., Promptify (GitHub, 2024c) for prompt engineering, and GPTCache (GitHub, 2024b) for reducing LLM API call expenses), post-processing (e.g., filtering out unwanted content), integration (e.g., integrating with websites, servers, or mobile apps), and plugin interactions (e.g., Zapier (GitHub, 2024d) for linking with external apps). On the other hand, businesses may prefer open-source LLMs to enable flexible customizations, which demand more sophisticated components, during stages such as pre-training, model reusing, fine-tuning (e.g., Labelbox (labelbox, 2024)), model conversion (e.g., Onnx (onnx, 2024)), quantization (e.g., TensorRT (nvidia GitHub, 2024)) and deployment. Behind the directly used components, there exists a multitude of sub-components that are transitively depended upon, further increasing the overall dependency complexity. Furthermore, the supplying types in the LLM supply chain are more varied than in traditional software supply chains. Data and models introduce additional supply types, including partial data reuse, model parameter reuse through data augmentation, fine-tuning, etc. Overall, leveraging LLMs is a systematic process that entails numerous components and supplying types across multiple stages, collectively referred to as the LLM supply chain.
Accordingly, the risks in the LLM supply chain have been significantly amplified due to an expanded attack surface and a growing number of participants. For stakeholders, it is essential to understand their providers and clients, how their artifacts are served, the risks associated with them, and the mitigation measures that are required. While the LLM supply chain shares similarities with the well-recognized open-source software supply chain (Ladisa et al., 2023), it also exhibits its uniqueness. For instance, attackers can exploit publicly accessible LLMs through prompt engineering to extract sensitive information (e.g., privacy leakage) or fine-tune open-source LLMs (eugeneyan, 2024) to embed malicious instructions within the model (e.g., backdoor attacks). Therefore, a comprehensive understanding of the types of components and supplying types in the LLM supply chain, thus to understand their corresponding risks is crucial to enable stakeholders to implement effective mitigation strategies.
While some literature touches on risks associated with the LLM supply chain, there is currently no paper that explicitly defines its scope, identifies inherent risks, and examines potential mitigation strategies in the LLM supply chain. Ladisa et al. (Ladisa et al., 2023) developed taxonomies for attack and defense techniques in software supply chains. However, their taxonomies do not account for LLM-specific components such as pre-trained LLMs, prompts, and data, making their survey less attractive for LLM supply chain stakeholders. Several surveys address the security and privacy challenges (Yao et al., 2024a; Cui et al., 2024b; Kumar et al., 2024; Yao et al., 2024a; Neel and Chang, 2023; Esmradi et al., 2023) and discuss the defense and mitigation strategies of LLMs. However, they did not discuss further what kinds of stakeholders and artifacts are affected, which fails to provide actionable insights and a clear scope to the corresponding stakeholders. Additionally, Wang et al. first (Wang et al., 2024b) outline a future research agenda for the LLM supply chain. However, they did not clearly define the scope of the elements that make up the LLM supply chain, making the covered challenges and opportunities that are less comprehensive than they could be. As the topic is still emerging and rapidly evolving, existing works may not comprehensively incorporate insightful ideas and perspectives from online preprints or web blogs. Therefore, we believe that a comprehensive LLM supply chain overview as well as a taxonomy of risks and clear guidance of mitigation in the LLM supply chain could be beneficial for both industrial and academic audiences. It could guide practitioners to pay attention to the LLM supply chain risks and provide a visionary map for academic researchers to devise new mitigation or defensive techniques.
In our paper, we provide a comprehensive overview of the LLM supply chain, detailing the stakeholders, composing artifacts, and the supplying types, as illustrated in Figure 1. We developed taxonomies of risk types, risky actions, and mitigations related to various LLM supply chain stakeholders and components. Specifically, we fist holistically collected both academic and online resources. Then, we summarize the risk scenarios as stakeholders performing risky actions on specific supply chain components, which in turn lead to distinct risk types. We illustrate the risk scenarios from the perspective of upstream contributors, downstream users, and administrators. Furthermore, we list the types of various mitigation measures in response to the LLM supply chain risks.
Our paper explores the technical and operational aspects of the LLM supply chain, offering a holistic understanding that draws valuable insights for researchers and engineers in software engineering, system architecture, software security, and data governance. The main contributions of our paper are as follows:
  • We presented a comprehensive overview of the LLM supply chain, including artifacts, stakeholders, and supplying types.
  • We develop a detailed taxonomy of risk and mitigations with regard to stakeholders, risk types and supply chain artifacts, providing actionable guidance for practitioners involved in the LLM supply chain.
  • We envision future challenges and opportunities in securing the LLM supply chain.
2. Large Language Model Supply Chain
We first provide our definition for the large language model supply chain. Then, we detail the stakeholders and composing components involved in the LLM supply chain, and provide a description of the LLM development stages. Lastly, we compare our work with existing surveys relevant to the LLM supply chain.
2.1. Definition
We present an overview of the large language model supply chain in Figure 1. While it may not capture the entire landscape exhaustively, we have aimed to make it as comprehensive as possible. The large language model (LLM) supply chain encompasses the ecosystem and sequence of processes integral to developing, training, deploying, and distributing applications that leverage large language models. This supply chain includes:
  • Stakeholders in various roles. Stakeholders take different roles based on the specific supply artifact and action. For instance, they may include developers contributing to the artifacts, organizations managing the platforms, or vendors providing cloud computing services.
  • Upstream artifacts in multiple forms. Upstream artifacts are those on which downstream applications have dependencies on. For example, these artifacts may include plaintext source code, binary executables, or serialized models.
  • Diverse supply relationships. As artifacts vary, the types of supply relationships differ. For instance, they may include augmentation relationships for data, cloning relationships for code, and merging relationships for models.
  • Development stages for various purposes. The development stages include a comprehensive scope for LLM application. e.g., data preprocessing, model pre-training, fine-tuning, integration, delivery, monitoring, and feedback.
2.2. Composing Components
We summarize five components in the LLM supply chain, i.e., stakeholder, artifact, toolchain, platform, and LLM applications (apps).
Figure 1. Overview of the Large Language Model Supply Chain
2.2.1. Stakeholders
We summarize four roles of stakeholders in the LLM supply chain, i.e., contributors, consumers, administrators, and user. Note that we categorize the stakeholders as roles because an individual can occupy a single role or assume multiple roles simultaneously, such as being both an contributor and an administrator in an open-source package platform.
Contributors. We collectively refer to contributors engaged in development activities and sharing artifacts, such as data, models, prompts, or third-party libraries, as contributors. Their roles vary depending on the types of artifacts they contribute.
  • Platform Contributors are responsible for contributing to hosting platforms (e.g., hosting artifacts), hardware platforms, and serving platforms (i.e., Google Cloud (Cloud, 2024), Amazon Web Services (Services, 2024)). They are obligated to provided secure, stable, isolated, and managed environments. i.e., malicious content detection, safe system images, authorized cloud accounts, etc.
  • Artifact Contributors collect, process, and upload various forms of data, code, prompts, and models to the hosting platforms. They can be data engineers, model developers, prompt engineers, and library developers.
  • Toolchain Contributors are responsible for developing, maintaining, and improving the tools used throughout the model lifecycle. e.g., Jupyter, TensorFlow for model development, Jenkins, GitHub Actions in CI/CD.They ensure that these tools are reliable, scalable, and compatible with the requirements from Toolchain Consumers.
Consumers. We distinguish consumers depending on their consuming components, including platform consumers, artifact consumers, toolchain consumers, and app consumers.
  • Platform Consumers are entities that use the platform infrastructure for various purposes, such as accessing hosting services, utilizing computational resources, or interacting with serving platforms for deploying and running models and applications.
  • Artifact Consumers are those who utilize the artifacts produced and shared within the platform. This includes data sets, code, prompts, and models.
  • Toolchain Consumers are individuals or teams that leverage the toolchain components for developing, deploying, and maintaining models. They use model development, CI/CD, and deployment tools to streamline the lifecycle of models and applications.
  • App Consumers are end-users or other systems that interact with applications built using the platform’s resources and artifacts.
Users. The users of LLM applications perform interactions with LLM applications in various forms. Their feedbacks (explicitly or implicitly) may be collected back to the LLM application.
Administrators. We summarize administrators as platform administrators and apps administrators.
  • Platform Administrator include two sub roles. The platform administrators manage and oversee the infrastructure, ensuring the system’s operational integrity, security, and performance. The project administrators, on the other hand, focus on managing repositories of artifacts, such as data repositories, model repositories, prompt repositories/websites, and library repositories. They hold the privileges to merge pull requests, manage versioning, and onboard new contributors.
  • Apps Administrator is responsible for managing the lifecycle of applications that rely on the artifacts and models within the platform. This includes deploying, configuring, and monitoring applications.
2.2.2. Artifacts
The LLM supply chain artifacts include data, model, prompts and code.
Data. Data serves as the foundational element in the workflow. It can be augmented and preprocessed from the original data to improve its quality, ensuring it is ready for use in training and fine-tuning the model. Retrieval methods bring in relevant data from various sources, facilitating the prompt generation process. They can be used as the data for pre-training or fine-tuning, or as the knowledge database for retrieval augmented generation.
Model. The model represents the large language model that is trained and fine-tuned using the processed data and instructions from the prompt. The model may be adjusted or modified by splitting, merging, serializing, or deserializing components as needed to optimize performance. It generates predictions or outputs based on the prompts used as the instructs.
Prompt. The prompt is crafted to instruct the model effectively, guiding it on what information to retrieve, process, or generate. Prompts can be refined to adapt to specific scenarios, leading to more accurate or tailored instructions that improve model outputs.
Code. The code encompasses the software and scripts utilized for data processing, model training, and prompt refinement. This includes code that directly invokes relevant library APIs, as well as the libraries themselves as dependencies. The code can be cloned for customization to enhance compatibility and functionality.
2.2.3. Platforms
The platforms are publicly accessible websites that provide essential infrastructure and services for distributing, and deploying various components of the LLM ecosystem.
Hosting Platform. These platforms facilitate the storage, sharing, and versioning of software packages, models, data, and related resources required for LLM development and deployment. e.g., Data Hosting Platforms (e.g., Kaggle (LLC, 2024) for datasets), Model Hosting Platforms (e.g., Hugging Face (Face, 2024) for pre-trained models), Repository Hosting Platforms (e.g., GitHub (GitHub, 2024a) for code repositories), Package Registries (e.g., PyPI (Foundation, 2024) for Python packages, Debian (Project, 2024) for Linux packages), Plugin Marketplaces (e.g., Visual Studio Marketplace (Microsoft, 2024) for development tools), and Prompt Sharing Platforms where users can share and discover prompt engineering techniques.
Hardware Platform. These platforms consist of specialized hardware designed to optimize the performance of LLMs. They include dedicated chips and accelerators (e.g., NVIDIA GPUs, Google TPUs) that provide the computational power necessary to handle large datasets and complex models, reducing training time and improving efficiency in deployment.
Serving Platform. These platforms offer cloud-based services to support the training, deployment, and maintenance of LLMs. They include Cloud Computing Platforms that provide scalable computing resources, such as Google Cloud (Cloud, 2024) and Amazon Web Services (Services, 2024). These platforms allow users to leverage high-performance infrastructure, including GPUs and TPUs, for intensive tasks like model training and inference.
2.2.4. Toolchain
We discuss the relevant toolchain with regard to model development toolchain, model deployment toolchain, and CI/CD toolchain.
Model Development Toolchain. Tools used for training or fine-tuning the LLMs. If an attacker implants malicious code in these tools, it could lead to corrupted models or abnormal outputs. Environment Image is the base environment (e.g., Docker containers) in which the model runs could be compromised, affecting system security or performance. Model Serializer and Deserializer save or load models. If compromised, they could modify the model’s internal state or introduce backdoors.
Model Deployment Toolchain. Tools specifically used for creating and deploying LLMs. Any manipulation or vulnerabilities here could compromise the model’s behavior or safety. Model Quantization Tool reduce the size of the deployment model, making it more efficient. If compromised, these tools could degrade model performance or introduce malicious behavior. Model Conversion Tool converts models between formats or versions. A poisoned conversion process could lead to abnormal model behavior or failures.
CI/CD Toolchain. Continuous Integration/Continuous Deployment systems that automate the software delivery process. A compromised CI/CD pipeline could introduce vulnerabilities during model updates.
2.2.5. Applications
Applications provide functionalities and services directly used by users. These applications include chatbots, automated customer service, text-generation tools, and more. Application developers integrate large models and API interfaces into real-world products to develop intelligent applications that meet specific needs. These applications directly serve end users, improving work efficiency or solving practical problems.
2.3. Model Development Lifecycle
Figure 2. Large Language Model Development Stages
Figure 2 illustrates the four essential stages of the large language model (LLM) lifecycle. Each stage includes specific processes crucial to building, deploying, and maintaining LLMs effectively. In the Preparation stage, foundational work is completed, including data pre-processing to clean and organize raw data, foundation model (FM) preparation to set up initial models, and prompt engineering to guide model outputs effectively. These steps ensure that the inputs are optimized for subsequent stages. The Training stage focuses on developing the model’s capabilities. Here, model pre-training introduces the model to large datasets, enabling it to understand language patterns. This is followed by fine-tuning on specific tasks to enhance accuracy and model compression to optimize the model for efficient deployment. In the Deployment stage, the trained model is integrated into real-world applications. This involves model integration to embed the LLM within applications, model conversion to make it compatible with different systems, and model serving to set up scalable infrastructure for user interactions. Finally, the Delivery stage ensures the ongoing performance and governance of the model. Model monitoring is conducted to track performance and detect issues, model maintenance includes regular updates and retraining to retain effectiveness, and model governance oversees the ethical and regulatory aspects, ensuring compliance and responsible usage.
2.4. Related Surveys
We conducted a survey of the security aspects of large language models (LLMs). Specifically, we gathered 12 surveys from our literature collection (see Section 3.1) that discuss various security-related topics, including LLM risks, privacy challenges, adversarial attacks, and defenses. For ease of comparison, we categorized the topics into two main areas: attacks and risks and defensive techniques. Additionally, we unified the terminology and definitions across the surveys to enable a clearer comparison.
Attack and Risks. In the context of attacks and risks, the Output Risk is extensively covered in eight survey papers, which refers to the potential for LLMs to generate harmful, untruthful, hallucinatory, or unhelpful content. Following this, Privacy Attacks are another major concern, with nine papers addressing various aspects such as membership inference attacks, privacy leakage (e.g., PPI) attacks, gradient leakage attacks, model inversion attacks, and attribute inference attacks. Besides, Prompts Attack is covered in 7 papers, including prompt injection attack and jailbreaking attack. Unfortunately, only a few papers (Pankajakshan et al., 2024; Cui et al., 2024b) mention Toolchain Attacks. Yao et al. (Yao et al., 2024a) discuss Supply Chain Vulnerabilities, and Wu et al. (Wu et al., 2024) highlight issues like malicious webtool misuse and cross-session access, which are closely related but account for a subset of Toolchain Attacks. Therefore, the topic of toolchain attacks was not explored in depth. Finally, only Neel et al. (Neel and Chang, 2023) addressed the issue of Copyright Risks.
Defenses and Mitigation. Defenses and Mitigation are suggested in accordance with the corresponding attacks and risks. For Prompt Attacks, existing works propose Input Sanitization (Pankajakshan et al., 2024; Iqbal et al., 2023). To mitigate Output Risk, approaches like Output Sanitization are recommended by several studies (Yao et al., 2024a; Cui et al., 2024b; Kumar et al., 2024). For Data Attacks, the literature suggests techniques such as Data Cleaning (Yao et al., 2024a), Data De-duplication (Neel and Chang, 2023), and Data Sanitization (Esmradi et al., 2023). However, these defensive techniques are not well-summarized across the surveyed works. While some methods, such as Differential Privacy, Federated Learning, or LLM Alignment, refer to specific technical measures designed to defend against particular attacks or mitigate associated risks, other defensive measures are less detailed. For instance, strategies like Output Detection and Input Sanitization are suggested to address Output Risk and Prompt Attacks, respectively, but they lack in-depth technical details on how these defenses are implemented.
Comparison to Existing Surveys. Our paper contributes the following contributions which are novel and incremental to existing surveys. First, we conducted a comprehensive survey in both academic literature and online resources regarding the open-source LLM supply chain. Second, we provide a comprehensive overview and highlight the key stakeholders and components in the LLM supply chain. Third, we summarize the risk scenarios as stakeholders performing risk actions on specific supply chain components, as well as risk types. We illustrate the risk scenarios from the perspective of three major stakeholders. Furthermore, we list the types of various mitigation measures in response to the LLM supply chain risks.
3. Methodology
In this section, we elaborate the methodology for our study. The overview of our methodology is presented in Figure 3.
3.1. Literature Collection
We focused our literature collection on academic literature and online resources. It helps to cover a comprehensive scope for our study. On the one hand, academic literature provides a rigorous understanding of the risks and mitigations in the LLM supply chain. On the other hand, online resources provide real-time, validated observations and practical examples that reflect current real-world practices.
3.1.1. Academic Literature Collection
The academic literature collection consists of two phases. First, we collected the initial papers set by querying with keywords that are central to our topic. Second, we expanded the paper set using the snowballing technique, incorporating both the references cited by the initial set and the papers that cite them.
Initial Paper Set. Two PhD students with over five years of experience in software supply chain management and two PhD students with over three years in LLM security have engaged in the process. Specifically, they first collected keywords from three domains. i.e., supply chain security (e.g., security, attack, malware, threat, vulnerabilities), LLM security (e.g., data poisoning, data provenance, model poisoning) and LLM DevOps (e.g., data validation, model validation, model evaluation). The remaining participants cross-checked the keywords, and together they compiled the final comprehensive set. Afterwards, we constructed queries based on the following rules using the advanced search interface on Google Scholar.
(LLM security AND supply chain security)
OR (LLM security AND LLM DevOps) OR
(supply chain security AND LLM DevOps)
OR LLM security
Figure 3. Approach Overview of our Literature Review
Furthermore, we excluded papers that do not align with the goal of our study. First, we filter out papers published before 2018. Second, we exclude short papers with fewer than 2 pages. Finally, we manually review the remaining papers to assess their relevance to the large language model supply chain. Consequently, this process leaves us with 127 papers.
Snowballed Paper Set. We identified cited and citing papers from the initial set, with each new group of relevant papers forming the basis for the next round of reviews. This iterative process continued until we reached a point where no new relevant papers were found. Ultimately, this process led us to a final set of 313 papers.
3.1.2. Online Resource Collection
In addition to collecting scientific literature, we have also gathered online resources which often include real-world attacks or the latest knowledge not yet covered by scientific literature.
Table 1 lists the websites where we collected online resources. We chose these websites for their wide coverage, reputation in the field of cybersecurity, professionalism, timely updates, and authoritative information. On the one hand, we leverage the same keywords used in the academic literature collection on the websites containing querying interfaces. e.g., Google. On the other hand, we manually inspect the websites to collect the relevant news or events.
Table 1. Websites of Online Resource
Websites Link
Google https://www.google.com/
MITRE ALTAS https://atlas.mitre.org/
GitHub Blog https://github.blog/
JFrog https://jfrog.com/blog
Sonatype https://www.sonatype.com/blog
Synk https://snyk.io/blog/
Tencent Security https://security.tencent.com/index.php/blog
Hacker News https://thehackernews.com/
Checkmarx https://checkmarx.com/blog/
Check Point https://blog.checkpoint.com/
BleepingComputer https://www.bleepingcomputer.com/news/security
The Register https://www.theregister.com/security
Security Intelligence https://securityintelligence.com/news/
heise https://securityintelligence.com/news/
SC Magazine https://www.scmagazine.com/
3.2. Literature Analysis and Taxonomy Construction
We manually inspected and categorized various elements of large language model risks and mitigations from the literature. First, we used open card sorting in our analysis. We began by thoroughly reviewing the literature set, and identified and highlighted terms, concepts, techniques, and strategies related to the large language model risks and mitigations.
We performed hierarchical tree structure construction, where we connected the identified categories into a structured, hierarchical framework that reflects the relationships between different risks and mitigation elements. We organize these categories into tree structures where the root represents the risks, stakeholders, and mitigations. To ensure accuracy and comprehensiveness, we refine our summarized structures according to the following criteria with multiple iterations.
  • Mutual Exclusivity. Assess whether each node is distinct from the others.
  • Accuracy. Verify the precision of node descriptions by consulting attack case studies and expert feedback, ensuring alignment with real-world attack patterns.
  • Hierarchy. Evaluate whether the tree’s structure is logical, ensuring broader attack categories at higher levels and specific methods at lower levels.
  • Comprehensibility. Test the clarity of node descriptions through expert panels or user feedback, ensuring that they are easy to understand.
  • Relevance. Ensure that each node reflects the latest risks and mitigations by reviewing recent research and industry standards (e.g., CVE, CWE, MITRE ATT&CK).
  • Coverage. Check whether the node includes all known attack techniques for its type, comparing against existing security frameworks and databases.
Each node in the structures is assessed by the four professionals independently. The structures are refined and re-evaluated iteratively until all professionals reach a final agreement.
4. Risks of Large Language Model Supply Chain
In this section, we summarize the risks in the LLM supply chain through illustrative risk scenarios. The scenarios are presented as risky stakeholders performing risky actions on specific supply chain components, which in turn lead to distinct risk types. We first discuss the risky stakeholders. Then, we study the risk types. Next, we propose the taxonomy of risky actions in the large language model supply chain. The risky stakeholders are detailed in the Section 2.2.1 and the supply chain components are covered in the Section 2.2. The risky actions are illustrated in Section 4.2 and the risk types are illustrated in Section 4.3. We present the risk scenarios regarding Risky Contributors, Risky Consumers, Risky Administrators and Risky Users in Figure 5, 6, 8 and 7, respectively.
4.1. Risky Stakeholders
We discuss the risky stakeholders in terms of risky contributors, risky consumers, risky administrators and risky users.
  • Risky Contributors.
    Camouflage Contributors. These contributors attempt to blend in with legitimate contributors, possibly hiding their true intentions.
    Attack Contributors. This group actively seeks to compromise the system by injecting malicious code, data, or models into the supply chain.
    Being Benign Contributors. Contributors in this category introduce vulnerabilities unintentionally, either through poor coding practices, lack of maintenance, or other factors that increase system risk.
  • Risky Consumers.
    Being Middleman. Entities that act as intermediaries, facilitating the exchange or integration of supply chain components between contributors and other consumers.
    Being Artifact Consumer. These stakeholders use artifact (e.g., data) for training, analytics, or other purposes.
    Being Platform Consumer. Entities that rely on platforms built with supply chain components, such as development or deployment environments.
  • Risky Administrators.
    Attack Administrators. Attackers can hack into an administrator’s account, causing severe consequences given the elevated privileges and control administrators hold over systems.
    Being benign administrators. Administrators who despite having good intentions, inadvertently introduce risks into the system through misconfigurations or other unintentional actions.
  • Risky Users. Users can be risky if their purposes are to exploit LLM application vulnerabilities, steal models, conduct jailbreaking or inference attacks, perform excessive access, or pollute the data using user feedbacks.
Figure 4. Risk Types
4.2. Risk Types
The LLM supply chain risks are classified into four main types. i.e., Security Risks, Privacy Risks, Delivery Risks, and Legal Risks. Figure 4 provides a comprehensive overview of the LLM supply chain risks across security, privacy, operational, and legal domains.
  • Security Risks (SR) include potential threats such as hacker use (SR1), system compromise (SR2), and the emergence of abnormal content (SR3) that could disrupt systems or compromise integrity.
  • Privacy Risks (PR) focus on data vulnerabilities, such as data leakage (PR1), personally identifiable information (PII) leakage (PR2), model stealing (PR3), code leakage (PR4), and prompt leakage (PR5), all of which can result in the unauthorized access or misuse of sensitive information.
  • Delivery Risks (DR) highlight issues related to the performance and maintenance of services, including performance degradation (DR1), system unmaintenance (DR2), and denial of service (DoS) (DR3) attacks that hinder service availability.
  • Legal Risks (LR) emphasize complications arising from copyright disputes, license disputes (LR1), and malicious use (LR2) of systems, which can lead to legal and compliance challenges.
4.3. Taxonomy of Risky Actions
We present the taxonomy of LLM supply chain risks with regard to the six composing elements in the LLM supply chain.
Figure 5. Risky Actions, Supply Chain Components and Risk Types Introduced by Risky Contributors
4.3.1. Risks on Third-party Libraries
Third-party libraries, including packages, frameworks, and plugins, have become essential at every stage of LLM application development. However, using these libraries comes with potential risks.
Implant Malicious Code. Attackers can develop malicious code and attempt to implant it through various methods, finally integrating it into the LLM supply chain. It can cause the risk of malicious use (LR2).
  • Dependency Confusion. A dependency confusion attack is a type of supply chain attack where an attacker publishes a malicious package to a public package repository with the same name as a popular private package. The package manager may unknowingly download the malicious package from the public repository instead of the intended private repository (PyTorch, 2022).
  • Typosquatting. Using typosquatting or combosquatting, attackers upload malicious packages with names that are nearly identical to popular legitimate packages. A minor typographical error during package installation can result in the inclusion of malicious code in the application. For example, PyPI has suffered from numerous malicious typosquatting packages. These malicious packages can lead to the leakage of Personally Identifiable Information (PII) and compromise the integrity and security of systems (CheckPoint, 2024). Rokon et al. (Rokon et al., 2020) identified over 7,500 instances of malware hosted on GitHub. Attackers may also employ more covert methods, such as spreading malware through forks (Cao and Dolan-Gavitt, 2022).
  • Compromising Legitimate Libraries. Attackers may target legitimate libraries already integrated into the LLM supply chain, leveraging their established presence and substantial user base. Attackers can conceal their malicious code within pull requests, subtly injecting harmful elements into the codebase. In addition, attackers may take control of the code from a previous maintainer and release new malicious versions (II, 2018).
  • Sharing Cloud Environments. Attackers can inject malicious code into cloud environments and make it publicly available to all cloud users. The rise in popularity of containers (e.g., Docker) has accelerated the trend of deploying applications in cloud environments. However, it in turn has become a hugely popular attack vector recently. The research team discovered that over a thousand Docker container images were found hiding malicious content (Chierici, 2022). Tomar et al. (Tomar et al., 2020) proposed a taxonomy of potential attacks at the container layer, revealing different types of threats that can arise within containerized environments.
4.3.2. Risks on Dataset
The datasets play a fundamental role, providing the extensive data required for model training, model refining and prompt generation.
Implant Poisoned Data. Attackers can create poisoned data and implant it through various methods, with the aim of having it used during critical stages of building the LLM supply chain. It can cause the risk of data leakage (PR1) and malicious use (LR2).
  • Data Poisoning. Attackers can manipulate the data for malicious purposes, and then upload it to public platforms (e.g., Kaggle) (Ma et al., 2019). Given that training large language models requires vast amounts of data, it is highly likely that these poisoned data sets will be used by LLM model developers. If the poisoned data is unintentionally leveraged into the LLM supply chain, it would result in the model performance degradation, particularly in its ability to generate the outputs accurately and efficiently (Shu et al., 2023; Wan et al., 2022; Shan et al., 2024). Alexander et al. (Wan et al., 2023) demonstrated that with only 100 poisoning examples, LLM produces consistent negative results or vulnerability output across hundreds of tasks.
  • RAG Poisoning. Retrieval-augmented generation (RAG) is widely incorporated in the LLM supply chain to provide the LLM with sufficient contextual knowledge. Similarly, the knowledge database can also be poisoned, where it can be manipulated the guide the LLM to generate specific, attacker-chosen responses to particular questions. Zou et al. (Zou et al., 2024) proposed PoisonedRAG, which specifically focuses on injecting poisoned texts into knowledge databases to steer language models to respond with predefined answers to questions selected by the attacker. Wang et al. introduce the concept of Poisoned-LangChain (Wang et al., 2024a), a novel indirect attack that exploits a poisoned external knowledge base.
4.3.3. Risks on Models
The open-source models provide accessibility, flexibility, and transparency, allowing developers to modify, customize, and integrate them into various projects.
Implant Model Backdoors. Attackers can manipulate publicly accessible models to alter their behavior for malicious purposes, ultimately affecting downstream users who rely on the compromised models. LLM may be composed of multiple components, including vocabularies, tokenizers, embedding layers, and auxiliary models (e.g., CLIP-ViT), etc (Liu et al., 2024; Radford et al., 2019; Chen et al., 2023). It can cause hacker use (SR1), system compromise (SR2), emergence of abnormal content (SR3), privacy risk (PR), data leakage (PR1), personally identifiable information (PII) leakage (PR2), model stealing (PR3), code leakage (PR4), and prompt leakage (PR5).
  • Component Dependency Attack. Attackers can manipulate the contents or structures of the composing models of LLMs, thereby conducting component dependency attacks. Huang et al.(Huang et al., 2023) firstly explore how manipulating the embedding dictionary using carefully designed rules could cause the model to produce specific outputs. Additionally, Cui et al.(Cui et al., 2024a) and Shayegani et al. (Shayegani et al., 2023) demonstrated that LLMs are not robust against attacks targeting the visual models used exclusively in multi-modal LLMs.
  • Backdoor attack. Backdoor attack embeds hidden backdoors in the model during the training process, allowing the compromised model to perform normally on benign samples but altered on the hidden backdoor input (Li et al., 2021; Chen et al., 2021; Pan et al., 2022). Yao et al. and Cai et al. (Cai et al., 2022; Yao et al., 2024b) studied the vulnerability of prompt learning algorithms commonly used in LLM to backdoor attacks.
Inference Attack. Inference attacks are types of privacy attack targeting data, where the responses of LLM models may inadvertently leak sensitive information, leading to unauthorized access, intellectual property theft, and privacy breaches. They include attribute inference attacks and membership inference attacks. Attribute inference attacks involve deducing sensitive information (e.g., race, gender, and sexual orientation) from the behavior or responses of LLM models, even if such information is not explicitly included in the training data. Robin et al. (Staab et al., 2023) and Pan et al. (Pan et al., 2020) conducted systematic studies on various advanced large language models, which have shown that LLMs can accurately infer sensitive data such as identity, genome, healthcare, and location information without any prior knowledge. Membership inference attacks aim to predict whether a data sample was included in the training data of LLM, undermining the trust between data providers and users. Mattern et al. (Mattern et al., 2023) explored the significant threat of membership inference in the real-world using neighborhood comparison. It can cause data leakage (PR1), personally identifiable information (PII) leakage (PR2), model stealing (PR3), code leakage (PR4), and prompt leakage (PR5).
Model Stealing. Model stealing refers to attacks where an adversary attempts to extract sensitive information about a machine learning model (e.g., model gradients, training data, and model parameters), thereby compromising the model’s privacy. It encompasses various model privacies, including gradient leakage (Gupta et al., 2022), training data extraction (Zhang et al., 2023b; Yang et al., 2023; Yu et al., 2023), and model information extraction (Carlini et al., 2024). It not only infringes the intellectual property rights of the model owner but also leads to security risks. For example, adversaries can create shadow models via model stealing, which are then used to generate adversarial examples, reverse-engineer sensitive data used in model training (Pan et al., 2020), or bypass any access restrictions or usage limitations in the original model (Shen et al., 2023). It can cause the privacy risk of model stealing (PR3).
Regarding LLMs, several model-stealing attacks have been reported to be effective (Zhang et al., 2023b; Yang et al., 2023; Parikh et al., 2022). Li et al. and Yu et al. (Carlini et al., 2021; Yu et al., 2023) conducted experiments to investigate the effectiveness of training data extraction on LLMs. Truong et al. (Truong et al., 2021) proposed techniques from the domain of data-free knowledge transfer to perform model extraction, successfully replicating the capabilities of the original model. Carlini et al. (Carlini et al., 2024) demonstrated precise and significant information extraction from black-box generative language models by recovering the embedding projection layer of a transformer model, given typical API access.
4.3.4. Risks on Prompts
Since LLMs are typically built from general knowledge, prompts are an effective way to incorporate domain-specific insights, directly enhancing the quality, relevance, and accuracy of the model’s output. The shared prompts allow a wider community to contribute and share specialized knowledge. However, shared prompts present several risks. In the context of shared prompts, those jailbreak prompts can be maliciously leveraged and pose challenges to the security mechanism of existing LLMs.
Implant Poisoned Prompts. Implanting poisoned prompts deliberately inserting harmful or misleading prompts into publicly accessible repositories or datasets. These poisoned prompts can be designed to manipulate the behavior of large language models (LLMs) by influencing their output in undesirable ways. Since shared prompts are widely shared and reused, such poisoned prompts can easily propagate through the supply chain, leading to inaccurate, biased, or even harmful model responses for downstream users who incorporate them into their applications. It can cause the performance degradation (DR1) of dilivery risks.
  • Prompt injection. Prompt injection is a type of attack that exploits malicious prompt words with trusted prompts against LLM applications. By exploiting the incapability of distinguishing “good” and “bad” prompts, prompt injection can cause malicious inputs to affect the model’s behavior and output, leading to information leakage or other security vulnerabilities. Stumpp et al.(Greshake et al., 2023) demonstrated the potential of prompt injection to indirectly execute arbitrary code within a Python interpreter in LLM applications, thereby gaining control over the host server (e.g., accessing environment variables, arbitrary file contents, executing DoS attacks, etc.). Chen et al.(Chen et al., 2024) and Shayegani et al. (Shayegani et al., 2023) further explored the threats of prompt injection using multi-modal inputs.
  • Jailbreaking. Jailbreaking attempts to subvert safety filters built into the LLMs themselves (Willison, 2024). Recent research (Wei et al., 2024; Deng et al., 2024; Zou et al., 2023) has used a variety of methods to make the LLMs LLMs responsive to restricted or insecure content (e.g., content that contains pornography, biased content, or assists the user in committing criminal acts). Wei et al. (Wei et al., 2024) investigates the reasons for the success of such attacks and how they occur, emphasizing the need for safety-capability parity. To assess the potential harm caused by jailbreaking, Shen et al. (Shen et al., 2023) devised the JailbreakHub framework including 1,405 jailbreak prompts.
Figure 6. Risky Actions, Supply Chain Components and Risk Types Introduced by Risky Consumers
4.3.5. Risks on LLM Applications
The LLM application performs various tasks using the LLM involving natural language processing, task-based dialogue, database search, etc. It can be a web application, an agent or a chatbot. The potential risks may be exposed in multiple ways.
Exploit LLM Application Vulnerabilities. The LLM Plugin Store (e.g., GPT store (OpenAI, 2024)) is a platform where users can discover, install, and manage plugins designed for enhancing the functionality of large language models. However, attackers can publish malicious plugins, guiding the model to output incorrect content through prompt injection (Wu et al., 2024). Moreover, some third-Party plugins could lead to account takeovers and data leakage (Newsroom, 2024; Kovacs, 2024a, b). Exploit lLM application vulnerabilities can potentially cause the hacker use (SR1), data leakage (PR1) and performance degradation (DR1).
Excessive Access. Users may overload the LLM application with excessive requests, potentially causing it to malfunction or become unresponsive. A Denial of Service (DoS) attack is a type of cyberattack that seeks to deplete computational resources, resulting in latency or rendering resources unavailable. Given that large language models (LLMs) demand substantial computational power, attackers can deliberately craft prompts to diminish the availability of these models (Derner et al., 2023). Shumailov et al. (Shumailov et al., 2021) demonstrated the feasibility of performing DoS attacks within the domain of LLMs. These attacks significantly increased the energy consumption and latency of the model (boosted to 10 to 200 times), raising researchers’ concerns about performance security in critical decision-making scenarios. Stumpp et al. (Stumpp, 2024) et al. used prompt injection to control the execution of dead-end code by the code executor of an LLM application, resulting in an unresponsive program host server. Excessive access can potentially cause hacker use (SR1), system compromise (SR2), emergence of abnormal content (SR3), data leakage (PR1), personally identifiable information (PII) leakage (PR2), model stealing (PR3), code leakage (PR4), and prompt leakage (PR5), and performance degradation (DR1).
Feedback Pollution. Existing LLM systems are often tweaked based on user feedback to improve their performance (Griffith et al., 2013). However, attackers may disrupt the performance of an LLM system by injecting a large amount of irrelevant data, leading to a surge in false feedback. Consequently, analysts at the victim organization may find themselves spending excessive time reviewing and rectifying erroneous inferences (MITRE, 2021). More seriously, LLM model developers may use this erroneous feedback during the instruction tuning phase to conduct reinforcement learning based on human feedback. This situation can not only lead to a decline in model performance but also cause the model to exhibit unpredictable behavior when facing real-world data (Wan et al., 2023; Baumgärtner et al., 2024). Sun et al. (Sun et al., 2024) demonstrate two proof-of-concept examples, stealthily manipulating a neural code generation system to generate code with vulnerabilities, attack payload, and malicious and spam messages. Feedback pollution can potentially cause the hacker use (SR1), system compromise (SR2), emergence of abnormal content (SR3), privacy risk (PR).
Perform Middleman Attack. A Man-in-the-Middle (MitM) attack involves an attacker intercepting communication between two parties without their knowledge. In the context of using the LLM Application or any chatbot platform, a MitM attack could occur if an attacker gains access to the communication channel, allowing them to alter or steal sensitive information being shared between a user and the AI system. Kang et al. (Kang et al., 2024) demonstrated that attackers may intercept data transmitted during a user’s interaction with LLM-based applications, leading to unauthorized access to sensitive information. Si et al. (Si et al., 2023) proposed that attackers can set up a fake API that mimics a legitimate one while offering a lower price to lure users. It not only allows the attacker to modify or steal sensitive data but also to exploit economic gains from the middle. The middleman attack can potentially cause data leakage (PR1), personally identifiable information (PII) leakage (PR2), model stealing (PR3), code leakage (PR4), and prompt leakage (PR5).
4.3.6. Risks on Platforms
Many platforms play important roles in the LLM supply chain, supporting various stages such as source code management, data hosting, model training, monitoring, and deployment. However, these platforms can introduce inherent risks.
Camouflage Contributors. Attackers can create fake accounts that imitate existing platform accounts (e.g., official accounts) to publish malicious data, code, models, and more. Users may be misled by these repositories, inadvertently using malicious tools to develop or deploy large models. It can cause the risk of malicious use (SR1).
Malicious Contributors. Attackers can obtain and exploit credentials of existing accounts or API tokens through various methods (News, 2018, 2022). An LLM red team organization (Team, 2024) has demonstrated in practice that once in possession of these credentials, attackers can gain unauthorized access to a range of resources and services associated with LLMs, potentially compromising the integrity, confidentiality, and availability of the targeted systems and data. Attackers can also use social engineering tactics to become maintainers of legitimate projects. Once they have obtained maintainer status, they can introduce malicious code or dependencies, which may go unnoticed by other contributors or users due to the perceived legitimacy of the new maintainer. For instance, in the XZ attack, the attacker gained maintainer access to the XZ project, embedding a backdoor in a critical dependency (Bals, 2024). It can cause the risk of malicious use (SR1).
Exploit Platform Vulnerabilities. Exploiting platform vulnerabilities refers to attackers identifying and taking advantage of weaknesses in the underlying infrastructure of an LLM or related application. These vulnerabilities can exist in the platform’s software, hardware, or configuration, allowing attackers to perform unauthorized actions such as gaining access to sensitive data, altering model behavior, or disrupting services. It can cause the risk of hacker use (SR1), data leakage (PR1) and model stealing (PR3).
  • CI/CD Flaws. Attackers can leverage flaws in AI cloud services to carry out attacks. Researchers from Wiz identified vulnerabilities in SAP’s AI Core service, which could have allowed attackers to access sensitive data of other tenants within SAP’s cloud infrastructure (scmagazine, 2024). By exploiting these flaws, attackers can retrieve AWS tokens, access AI training data, and potentially cause data theft, service disruption, and even supply chain attacks within LLM applications.
  • Privacy Leakage. Attackers can exploit secrets found in public artifacts generated by automated CI/CD pipelines. Recent research revealed that artifacts produced by GitHub Actions workflows may contain sensitive tokens, such as GITHUB_TOKEN and ACTIONS_RUNTIME_TOKEN, which attackers could use to inject malicious code (French, 2024). When users download and utilize these compromised artifacts, it can result in a supply chain attack.
Figure 7. Risky Actions, Supply Chain Components and Risk Types Introduced by Risky Users
4.3.7. General Risks to LLM Supply Chain Component
Apart from risks that are specific to the six composing elements, we elaborate on general risks to LLM supply chain components.
Introduce Risky Components. A benign contributor may not realize that the components they introduce have hidden vulnerabilities or security flaws. These risky components can be exploited by malicious actors once deployed. For example, an LLM application might rely on a library that appears secure but contains a hidden vulnerability. When the benign contributor integrates this library, they inadvertently expose the LLM and its users to potential exploitation. Attackers can leverage these risky components for data theft, model manipulation, or unauthorized access. It can cause the risk of hacker use (SR1), data leakage (PR1) and model stealing (PR3).
Discontinue Maintenance. When a contributor stops maintaining a component, updates for security patches, bug fixes, and compatibility improvements are halted. It can lead to vulnerabilities if the abandoned component contains outdated dependencies that are prone to exploitation. Moreover, the discontinuation can disrupt further development by creating compatibility issues, reducing model performance, or complicating integration with newer systems. Furthermore, it poses significant risks because malicious actors could take advantage of this by forking the abandoned project and inserting malicious code. It can cause the risk of unmaintained (DR2).
Release Vulnerable Code. The benign contributors may submit code that contains security flaws or exploits without malicious intent. Such vulnerabilities could range from weak encryption practices, insecure data handling, or exploitable logic errors. While the contributor’s intent is not harmful, the inadvertent introduction of security issues can still have serious consequences, potentially leading to supply chain attacks. It can cause hacker use (SR1), system compromise (SR2), and the emergence of abnormal content (SR3).
Raise License Disputes. While open-source licenses generally grant permission to use, modify, and distribute the code, the owners retain the copyright. It introduces potential risks for users who download and use the code, model or data if contributors claim copyrights on their code. They may enforce restrictions or take legal action if users violate the terms of the license. In 2020, CoKinetic Systems sued Panasonic for allegedly violating the GPLv2 license, with the lawsuit seeking financial compensation (Murray, 2020). It can cause the risk of license disputes (LR1).
Figure 8. Risky Actions, Supply Chain Components and Risk Types Introduced by Risky Administrators
5. Mitigation of Risks in Large Language Model Supply Chain
We elaborate on the mitigation of risks in the LLM supply chain with regard to the seven aspects.
5.1. Mitigation to Risks from Third-party Libraries
Malware Detection. Malware detection identifies viruses, worms, trojans, and backdoors from third-party libraries in both binaries and source code. Key techniques include signature comparison, static analysis, and dynamic analysis. Signature-based detection tools, like VirusTotal (VirusTotal, 2023), are effective for identifying binary malware. To detect malware within package registries, static and dynamic analyses are widely applied. Recent research has advanced these approaches by modeling the behavior of malicious packages using static characteristics alone (Huang et al., 2024b; Zhang et al., 2023a), as well as integrating both static and dynamic analyses to enhance detection accuracy (Duan et al., 2020; Huang et al., 2024a). This measure can mitigate the risk SR2.
Software Bill-of-Materials Analysis. A Software Bill of Materials (SBOM) is a list of inventory components included in imported third-party libraries. SBOM analysis provides visibility into all third-party libraries and their versions, helping users avoid unvetted or potentially risky libraries. Several commercial tools offer SBOM analysis, including Sonatype Management (Sonatype, 2008), BlackDuck (BlackDuck, 2002), and Microsoft’s SBOM tool (Microsoft, 2022). By utilizing SBOM analysis, users can enhance security and improve overall software quality throughout the development lifecycle. This measure can mitigate the risk SR2 and SR3.
5.2. Mitigation to Risks from Dataset
Data-Deduplication. In the face of privacy risks, we can leverage data deduplication on the dataset. Jagielski et al. discovered that model privacy attacks are more effective on the training dataset that appears multiple times (Jagielski et al., 2023). Removing duplicate text from the training data  (Abadi et al., 2024) helps protect against model memorization of sensitive information, making privacy attacks less effective. This measure can mitigate the risk PR2.
Data Encryption. Data encryption protects data confidentiality by converting it into an unreadable format, consisting of three types: symmetric encryption (Huang et al., 2020), asymmetric encryption (Lin et al., 2023; Huang et al., 2022), and hybrid encryption (AbdElnapi et al., 2016). With effective key management mechanisms, the data cannot be deciphered, significantly reducing the risk of data privacy breaches. This measure can mitigate the risk PR1 an PR2.
Data Sanitization. Data sanitization cleans input data to defend against data poisoning and backdoor attacks by applying techniques like data denoising, filtering, and smoothing, which effectively remove adversarial noise while retaining legitimate information  (Xu et al., 2017). Additionally, these methods help prevent data leakage by ensuring that shared or processed data cannot be traced back to individuals. This measure can mitigate the risk SR3 and PR1.
Data Watermark. Data watermark protects data privacy by embedding specific watermark identifiers (Sun et al., 2023; Tang et al., 2023) into the training dataset. During the training process, the model is trained using watermarked data, ensuring that it generates output text containing the corresponding watermark when given specific inputs, thus achieving the goal of identifying and tracing the data source. This measure can mitigate the risk PR1.
Differential Privacy. Differential privacy (Bagdasaryan et al., 2020) protects the private information of individual records in a dataset (Dwork et al., 2006, 2014). Specifically, a differential privacy mechanism adds noise to each training data point. It ensures that even if a particular data point is replaced or removed, the model’s output will not be significantly affected, which effectively prevents attackers from inferring private information in the training data by analyzing the model’s outputs. This measure can mitigate the risk PR2.
5.3. Mitigation to Model Risks
Model Signature. Model signature helps to track and manage the origin of models. Jiang et al. (Jiang et al., 2023) described useful attributes for representing the model signature, such as provenance, reproducibility, and portability. Recording and verifying key attributes, provides a reliable reference point, allowing researchers or engineers to trace and verify the model’s training process, thereby determining whether the model has been maliciously modified or compromised. This measure can mitigate the risk PR3.
Model Watermark. Model watermark embeds traceable information into the text generated by LLMs to verify the model’s source and authenticity without degrading output quality. Techniques such as logits modification (Kirchenbauer et al., 2023), entropy-based embedding (Lee et al., 2023), multi-bit payload (Wang et al., 2023), and token sampling (Kuditipudi et al., 2023) enable efficient watermarking by embedding verifiable identifiers that protect against model-stealing, safeguarding intellectual property and ensuring model integrity. This measure can mitigate the risk PR3.
Model Obfuscation. Model obfuscation protects the security and privacy of the LLMs. It modifies the model’s structure or parameters to prevent attackers from being reverse-engineered or maliciously exploited. Zhou et al. (Zhou et al., 2023) proposed an active defense solution for model theft via automated weight obfuscation. This measure can mitigate the risk PR3.
Model Alignment. Model alignment ensures that a model’s outputs and behaviors align with human values and expectations. Reinforcement Learning from Human Feedback (Ouyang et al., 2022) incorporates human feedback into the model’s training process, guiding the model to learn which behaviors are appropriate or not. This measure can mitigate the risk SR3.
Adversarial Training. Adversarial training is a machine learning technique designed to enhance model robustness by exposing the model to adversarial examples (Xhonneux et al., 2024) during training. By learning to recognize and handle these malicious inputs, the model becomes more resilient to attacks and unpredictable behavior. This measure can mitigate the risk SR3.
Federated Learning. Federated Learning is a decentralized approach where multiple clients collaboratively train a shared model without transferring their individual data to a central server. It prevents the aggregation of malicious updates by identifying and excluding anomalies in the distribution using unsupervised or supervised machine learning techniques  (Li et al., 2020; Zhang et al., 2022). By leveraging three-stage defenses (pre-aggregation defense (Pre-AD), in-aggregation defense (In-AD), and post-aggregation defense (Post-AD)), federated learning ensures that the aggregated model remains secure and resistant to backdoor attacks or malicious data manipulation, while still benefiting from decentralized data training. This measure can mitigate the risk SR3.
5.4. Mitigation to Prompt Risks
Prompt Sanitization. Prompt sanitization filters, formats, or regularized prompts are provided to LLMs to ensure that the model does not misinterpret their original intent, thereby preventing abnormal content or malicious behavior. (Suo, 2024; Robey et al., 2023) have been shown to effectively prevent jailbreak in LLMs. Suo et al. (Suo, 2024) applied to sign to sensitive elements (e.g.,’delete file’) in the user’s prompt. Robey et al. (Robey et al., 2023) randomly perturb characters in a given input prompt (e.g., inserting random characters). Chen et al .(Chen et al., 2024) separates prompts and user data into two independent channels to prevent prompt injection. This measure can mitigate the risk SR3.
Robust Tuning. Robust tuning enhances the defense of LLMs against prompt risks by incorporating specific techniques during tuning. For example, Yue et al. (Yue et al., 2022) fine-tuning a generative language model with differential privacy to generate privacy-preserving synthetic text. Ozdayi et al. (Ozdayi et al., 2023) leverage prompt-tuning to control the extraction rates of memorized content, avoiding attack modifying LLM weights. This measure can mitigate the risk SR3.
Model Structure Optimization. Model structure optimization aims to enhance the robustness of LLMs’ inference by modifying their structure to defend against malicious prompts. Chen et al (Chen et al., 2022). Employ homomorphic encryption to enable privacy-preserving inference for Transformer-based models. Li et al. (Li et al., 2022) achieve efficient and private Transformer inference using Secure Multi-Party Computation (MPC). This measure can mitigate the risk PR1.
5.5. Mitigation on LLM Applications Risks
Secure Auditing and Permission Management. Secure auditing and permission management are crucial for mitigating risks associated with LLM plugins. This approach includes enforcing strict vetting processes to ensure that only trusted plugins are integrated into the system, implementing anomaly detection systems to identify unusual behaviors indicative of malicious activity, and conducting regular security audits to assess plugin security and compliance. Additionally, employing role-based access control (RBAC) limits permissions, ensuring that users and plugins have only the necessary access to perform their functions, thereby reducing the attack surface. By maintaining a comprehensive logging system, organizations can track interactions and quickly respond to potential security incidents, fostering a proactive security posture. This measure can mitigate the risks SR1, SR3 and LR2.
Authorized and Restricted Access. Authorized and restricted access is a key defense mechanism against several risks in LLM applications. For feedback pollution, restricting access to authorized users ensures that only credible feedback is collected for model training. For Middleman (MitM) attacks, by enforcing strong access controls, such as mutual authentication and encryption, only authorized clients can communicate with the LLM. For excessive access, limiting access to verified users via rate-limiting mechanisms mitigates the risk of excessive API requests that can overload the system, leading to Denial of Service (DoS) attacks. By ensuring only authorized users can access the LLM, and controlling the frequency and scope of interactions, systems can prevent abuse and maintain availability. This measure can mitigate the risks DR1, DR2 and DR3.
5.6. Mitigation to Platform Risks
Anomaly Monitoring. Platforms can apply authorized and restricted access for untrusted accounts. e.g., multi-factor authorization (PyPI, 2023). Apart from that, they can adopt anomaly monitoring by observing developing activities to identify any suspicious behavior, such as unauthorized access attempts, changes to account settings, or unexpected pull requests. For example, Danielle et al. (Gonzalez et al., 2021) leveraged commit messages to detect anomalies and malicious commits. This measure can mitigate the risk SR2.
Mitigating Platform Vulnerabilities. Platforms should also ensure that their CI/CD systems are following best security practices, such as least privilege, secure credential storage, and network segmentation, keeping all CI/CD tools and dependencies up to date with the latest security patches to protect against known vulnerabilities (scmagazine, 2024), and regular security assessments and vulnerability scans of the CI/CD environments to promptly identify security flaws. This measure can mitigate the risk SR2.
Session Isolation and Privacy Protection. Using sandboxing in platforms isolates user sessions within separate environments, preventing cross-session attacks. By restricting access to shared resources and data, sandboxing ensures that malicious activities in one session cannot affect others. This containment helps protect sensitive information and enhances overall security by limiting the impact of potential breaches. Besides, platforms can utilize secrets management tools, log scanning, and sanitization to ensure that sensitive information like tokens and API keys are not exposed in repositories or logs (French, 2024). This measure can mitigate the risks SR2, PR1, PR2, PR3 and PR4.
5.7. Mitigation to General LLM Supply Chain Component Risks
Enhancing Security and Legal Risk Awareness. Developers and organizations should be vigilant when incorporating components into their software systems. On the one hand, they should actively monitor for known vulnerabilities in third-party libraries and frameworks, thus reducing their exposure to potential threats. On the other hand, they should familiarize themselves with licensing agreements, intellectual property rights, and the potential liabilities tied to both open-source and proprietary components. This measure can mitigate the risks SR2 and LR1.
Training for Secure Coding. Adhering to secure coding practices is fundamental to minimizing software vulnerabilities. This includes following best practices for input validation, implementing robust error handling, and adhering to established secure coding guidelines. By prioritizing these practices, developers can substantially lower the likelihood of introducing security flaws into their applications, thereby safeguarding user data and assets. This measure can mitigate the risk SR2.
Preparing Alternative Components. To mitigate the risks associated with reliance on components that may become unsupported or discontinued, it is important to seek alternative solutions. Moreover, advanced replacement techniques are also needed to help migrate the original component into a new one with the least effort and cost. By diversifying their dependencies and preparing for potential discontinuations, the stability and security of the LLM-driven application can be enhanced. This measure can mitigate the risks SR2, DR2 and DR3.
6. Discussion
We discuss future challenges and opportunities towards mitigating the risks in the LLM supply chain.
6.1. Future Challenges
Comprehensive and Precise Decomposition of the LLM Supply Chain. Decomposing the LLM supply chain is essential in understanding and mitigating risks in the LLM-related ecosystem, ensuring compliance, and enhancing transparency across the LLM lifecycle. First, mapping an entire LLM supply chain is a complex task. As the scope and applications of LLMs expand, defining clear boundaries for the LLM supply chain becomes nearly impossible. Nevertheless, identifying a clear scope of relevant components is crucial to enable a comprehensive analysis of the LLM supply chain, including its core artifacts, stakeholders, and supply relationships. This clarity allows each stakeholder to understand their role and responsibilities within the supply chain effectively. Second, the unique dependencies of the LLM supply chain adds further complexity. Unlike traditional software supply chains, the LLM supply chain includes special dependencies—such as models embedding external knowledge (e.g., data or code). The models do not directly use external resources. Instead, they encode knowledge into tensors and weights, obscuring the origins and composition of these elements. Third, LLM artifacts (e.g., models, datasets, and prompts) differ fundamentally from software artifact, making it difficult to directly apply standard version management. This limitation complicates efforts to decompose the LLM supply chain effectively and hinders accurate tracking and management.
Compliance with Government Laws and Guidelines. The growth of new LLM applications is expected to accelerate with the increasing availability of commercial and open-source LLMs, leading to their adoption across a wide range of industries from different countries. Since industries across countries may be subject to varying regulations and requirements, ensuring compliance with government laws and guidelines is essential. First, as LLMs often rely on data sourced from various online platforms, verifying the legality of training datasets to see if it is intellectual property (IP) free is essential. Given that the conclusion of dataset composition may be indirect, it is challenging to establish a common sense for copyright infringement or IP violations solely based on the model’s generated output or trained weights. Furthermore, the providers may deliberately wipe traces of the IP for their benefit. Therefore, it is also crucial to fight against those actions. Second, the vast data required to train LLMs includes sensitive information that may pose privacy risks. Ensuring compliance with data privacy laws like GDPR, alongside securing data through encryption and anonymization is also a critical issue. Third, LLMs are prone to inheriting biases present in their training data, which can persist in the downstream application, affecting their performance and trustworthiness. Developing robust strategies for identifying and mitigating biases is crucial for ethical AI.
Vigilant Prevention of Adversarial Attacks. On one hand, with reliance on third-party data sources, libraries, proprietary frameworks, and cloud providers, there is a risk of supply chain vulnerabilities and limited control over model updates. This creates security concerns as well as challenges in maintaining compatibility and transparency. On the other hand, LLMs are increasingly targeted by adversarial attacks, such as data poisoning or model inversion, which can compromise data integrity or privacy. Developing robust defenses will be necessary to ensure secure usage.
6.2. Opportunities
Collaborative Open-Source Ecosystem. The prosperity of LLMs is inherently linked to the open-source community. The techniques for effectively leveraging the open-source ecosystem around LLMs are still underdeveloped. Open-source contributions should incorporate collaborative risk assessments and secure coding practices to prevent vulnerabilities. Additionally, establishing ethical guidelines for contributions, such as promoting diversity in training data and addressing bias, can ensure that open-source models are developed responsibly. Furthermore, there is a need for shared benchmarks and evaluation standards that enable developers to compare models transparently, fostering objective performance improvements and driving innovation.
Enhanced Transparency and Traceability. Improved transparency in sourcing especially in data and models during the whole LLM application developing process can build user trust, ensure accountability in LLM development, and foster ethical AI practices. There are three potential approaches. First, creating detailed audit trails for data usage and transformations in training pipelines can enhance accountability and support error tracking. Second, allowing controlled access to different data and model components supports transparency while protecting sensitive information. Third, tracking changes across model versions, from initial models to refined iterations, enables stakeholders to understand the model’s evolution and track back any issues related to specific changes.
Regulatory and Standards Development. Establishing standardized practices and regulatory frameworks for the LLM supply chain can help align practices across different providers, ensuring security, compliance, and interoperability. On the one hand, we need to define clear guidelines on data handling, including data anonymization and user consent, which can help prevent privacy infringements. On the other hand, regulatory frameworks could include guidelines for reducing bias, ensuring fairness, etc.
Improved Defense Mechanisms. Developing stronger defenses against adversarial threats can enhance the resilience of LLMs against cyber threats and tampering. Three potential approaches can be considered. First, implementing systems to detect unusual activity or output patterns can help identify and mitigate adversarial attacks effectively. Second, utilizing secure delivery channels from an identified contributor can prevent untrusted modifications to LLMs. Third, establishing automated models or library updates can address vulnerabilities in LLM applications promptly, reducing the risk of exploitation.
7. Conclusions
The rapid growth of large language models has transformed numerous industries, creating an intricate supply chain that may incur potential risks. This paper presents a comprehensive overview of the LLM supply chain. We identify and categorize the risks inherent in this supply chain, framing them through stakeholders, risky actions, risk types, and specific supply chain components. Additionally, we provide a taxonomy of mitigation strategies, offering actionable guidance for stakeholders seeking to navigate and secure the LLM supply chain. We also highlight emerging challenges and opportunities in securing the LLM supply chain, aiming to inspire further research into robust defenses and proactive security measures.